OpenStudy (anonymous):

Someone please help, I Will medal! What is the quotient of (3+i)(3-i)/-(2+2i)^2 expressed in simplest form?

OpenStudy (anonymous):

(9-i^2)/(4i^2+8i+4)=10/8i=5/4i=-5/4i